The electric field at (30,30) cm due to a charge of -8nC at the origin in N/C
 

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Distance between origin and point (30,30)cm or (0.3,0.3)m isd=sqrt{0.3^2+0.3^2}=0.3\sqrt{2}m

Therefore, electric field at the point=\dfrac{1}{4\pi\epsilon_0}\dfrac{q}{d^2}=9\times 10^9\times \dfrac{8\times 10^{-9}}{0.18}N=400N towards the origin.